Find the payment that should be used for the annuity due whose future value is given. Assume that the compounding period is the same as the payment period.

​$21,000​; quarterly payments for 16 years; interest rate 5.5​%
Answer #1

solution or P= 21,000 8=5.5% T 16 A = 21000(1+. 5.5 100 T = 16 years H=p (1+2) ° (1+ (A = 49460.51669) Total amount paid a= $